OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 6 inches and the radius is 7 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi. one hundred thirty-one point eight eight square inches two hundred sixty-three point seven six square inches seven hundred ninety-one point two eight square inches nine hundred twenty-three point one six square inches

OpenStudy (jordanloveangel):

(I believe the word you're looking for is the Greek letter "pi", not the dessert frequently served with ice cream.) Area of a cylinder = πDh = 2πRh = 2(3.14)(6)(8) = 301.44 square inches.

OpenStudy (jordanloveangel):

SA=2pi*r*height SA=3.14*(2*6)*8= 301.44 inches squared

OpenStudy (anonymous):

lol I know um is it A

OpenStudy (jordanloveangel):

yes

OpenStudy (jordanloveangel):

aye u o me 100 medal

OpenStudy (anonymous):

yessssssssssssssssssss thank u I have to count slow so just count 100 medals

OpenStudy (jordanloveangel):

lol its al right now need guys give me 1 its okay im messing around

OpenStudy (anonymous):

;) have a blessed day!